It is found out that 13% of students enrolled in College Algebra at a large university fail the first time they take it. A psychology professor claims that this could be reduced with a program of counseling, and so on. A random sample of 850 students enrolled in College Algebra take this program and 99 fail anyway. Is this a significant improvement at the 0.05 level?


1
Expert's answer
2021-06-08T04:27:13-0400

Hypothesized Population Proportion "p_0=0.13"

Favorable Cases "X=99"

Sample Size "n=850"

Sample Proportion "\hat{p}=\dfrac{99}{850}\approx0.11647059"

Significance Level "\alpha=0.05"

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p\geq0.13"

"H_1:p<0.13"

This corresponds to a left-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a left-tailed test is "z_c=-1.6449."

The rejection region for this left-tailed test is "R\{z:z<-1.6449\}"

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}"

"=\dfrac{\dfrac{99}{850}-0.13}{\sqrt{\dfrac{0.13(1-0.13)}{850}}}\approx-1.1729"

Since it is observed that "z=-1.1729>-1.6445=z_c," it is then concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is less than "0.13," at the "\alpha=0.05" significance level.

Using the P-value approach:

The p-value is "p=P(Z<-1.1729)=0.120418," and since "p=0.0.120418>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is less than "0.13," at the "\alpha=0.05" significance level.


Therefore, there is not enough evidence to claim that this is a significant improvement at the 0.05 level.